OpenStudy (anonymous):

A recipe calls for 4 L of water. Arlene wants to make 0.3 more servings than the recipe makes. Which expression represents how many liters of water Arlene should use? A. 3/10 x 4 B. 13/10 x 4 C. 7/10 x 4 D. 17/10 x 4

OpenStudy (some.random.cool.kid):

maybe A

OpenStudy (anonymous):

It is probably C. i think it is

OpenStudy (anonymous):

Well, if I were to find how many more liters I would need, I would do the following:\[4/10=.4\] This is 1/10th of the initial serving amount or .1, but I need .3 so I'm going to multiply that by 3. \[.4 * 3=1.2\] This is how much more I need. So I add the initial amount to this amount:\[4+1.2=5.2\] B is equal to 5.2. Though I am unfamiliar with the methods used in the question you provided. But yes, B is the answer.